# Atmospheric stability sets maximum moist heat and convection in the midlatitudes

### v3

@@ Abstract
Extreme moist heatwaves pose a serious threat to society and human health. To manage heat-related risks, it is crucial to improve our understanding of what limits the maximum near-surface moist heat. This question has been extensively studied for moist heat in the tropics under the assumption of moist neutrality, but it remains less explored over midlatitude continents, where this assumption is not applicable. The frequent concurrent occurrence of extreme moist heat and intense convection over midlatitude land reinforces the prevailing hypothesis that deep convection terminates heatwaves and sets their peak intensity. Here, a theory is developed based on the onset of convection to provide a theoretical prediction for both the maximum near-surface moist heat and potential convection over midlatitude land. Our framework demonstrates that the intensities of moist heat and potential convection are constrained by a preexisting energy barrier in the lower free troposphere, which suppresses convection and enables the buildup of moist heat and convective instability. This barrier, often marked by a temperature or energy inversion, varies only slightly during moist heatwave evolution. Our results suggest the potential to predict maximum moist heat and convection from predefined lower free tropospheric properties and offer insights into the evolution of extreme heat and convection across climate states.
This manuscript was compiled on January 22, 2025
@@
Moist heatwaves, which combine the effects of high temperature and humidity, pose significant risks to public health and societal outcomes (1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12). While previous research has primarily focused on the dynamics and changes of moist heat in tropical and subtropical regions (13, 14, 15, 16, 17, 18, 19), the drivers and characteristics of extreme moist heat over midlatitude continents remain poorly understood. As climate change is driving severe moist heat risks poleward into the midlatitudes (20, 21, 22, 23, 24, 25, 26, 27), advancing our theoretical understanding of the limits of maximum moist heat in these regions has become increasingly urgent. Our goal in this study is to determine what sets the maximum moist heat over midlatitude land.
@@
Near-surface wet bulb temperature (WBTs) is a key measure of moist heat stress, as it combines the influence of temperature and humidity (28, 3, 29, 4, 23, 16, 26, 30). By definition, WBT represents the temperature that an air parcel would attain if cooled adiabatically to saturation at constant pressure by evaporating water into it (28, 31, 32). During this adiabatic process, the air parcel conserves its moist static energy (MSE) or the equivalent potential temperature, and thus WBTs can be directly derived assuming the conservation of near-surface MSE (MSEs) (33, 16, 15, 30, 34). Unlike the non-linear WBT formula, MSE is a linear function of temperature, specific humidity, and geopotential height, making it simple to use (15, 35). Accordingly, this study employs MSEs, defined as the sum of sensible heat, latent heat, and geopotential energy at 2 meters above the ground, to measure near-surface moist heat. Overall, MSEs reasonably represents moist heat and, similar to the annual maximum WBTs (Fig. 1), the annual maximum MSEs recovers well the spatial distribution of extreme moist heat over midlatitude land (Fig. 1) and aligns closely with WBTs (Fig. S1; detailed in ). In this work, we use the 3-hourly ERA5 reanalysis data (36, 37) during 1980–2022 for land grids between 35∘N and 75∘N with elevation lower than 1000 m (detailed in ). Our analysis focuses on the annual maximum MSEs to explore the potential constraint on the maximum moist heat over midlatitude land.
@@
Previous studies have commonly assumed a strong coupling between the near-surface atmosphere and the free troposphere during extreme heat through deep moist convection, such that MSEs is constrained by free troposheric saturated moist static energy (MSE∗) given the state of convective quasi-equilibrium (38, 39, 40, 3, 41, 42, 43, 44, 16, 14, 4). Under the assumption of a moist-neutral atmospheric column, this framework has been widely used to study extreme heat in the tropics (14, 16, 18) or the global mean state (4). With the neutrality assumption, a recent study (45) has applied this framework to investigate extreme dry heat over midlatitude continents, suggesting that MSEs is limited by 500-hPa MSE∗ (i.e., MSEMSE), thereby setting an upper bound for maximum surface air temperature. We first test the validity of this theory by comparing MSEs with MSE during the annual maximum moist heat over midlatitude land. Our analysis reveals that the theory does not hold for extreme moist heat, as MSEs is consistently and significantly greater than MSE in nearly all cases (more than 96% of the midlatitude land grid points; Fig. 1–). Furthermore, the distribution of MSE is largely zonally uniform (Fig. 1), which does not capture the strong zonal variation observed in the MSEs pattern (Fig. 1).
@@
These results indicate a violation of the assumption of moist neutrality because, unlike quasi-equilibrium convection in the tropics, severe convection over midlatitude continents involves the strong storage of potential energy over a short time period (46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56), which potentially interacts with surface heat in distinct ways. Here, we propose a new theoretical framework that accounts for the onset of convection, which enables surface heat and convective instability to evolve beyond the neutral condition, and ultimately provides a tight constraint on both near-surface moist heat and potential convection over midlatitude land.
@@
We first demonstrate that deep convection occurs concurrently with the annual maximum moist heat (Fig. 2), supporting the hypothesis that convective instability acts to terminate heatwaves and constrain maximum near-surface heat (4, 45, 57, 58, 17). The potential intensity of convection is measured by convective available potential energy (CAPE; detailed in ), with values of 1000 J kg-1 or higher typically deemed sufficient to support strong to severe convection, according to the National Weather Service guidelines (59). We define CAPE at the time of the annual maximum MSEs as critical CAPE (denoted CAPEc), which measures the potential convection that terminates the annual maximum moist heat. The CAPEc over most midlatitude land regions exceeds 1000 J kg-1, with values ranging from 3000 to 4000 J kg-1 across western Europe extending to Northeast China, and reaching 6000 J kg-1 or higher over eastern China and central North America (Fig. 2). The high values of CAPEc represent substantial convective instability in the atmosphere during extreme moist heat, indicating a significant departure from the neutrality assumption. Meanwhile, critical convective inhibition (denoted CINc; detailed in ), a measure of the energy barrier in the lower free troposphere that suppresses the release of CAPE, is sufficiently low (generally less than 25 J kg-1) to allow the initiation of deep free convection in the midlatitudes (Fig. S2).
In fact, the deep convection that terminates the maximum moist heat is likely the most intense convection of the year across many midlatitude land areas, as CAPEc (Fig. 2) aligns closely with the annual maximum CAPE (Fig. S2). This alignment is particularly evident in Europe and central North America, where the difference between CAPEc and the annual maximum CAPE is less than 250 J kg-1 (Fig. S2). In these regions, over 50% of the years during 1980–2022 feature simultaneous annual maxima of CAPE and MSEs (Fig. S2), emphasizing the frequent co-occurrence of extreme moist heat and intense convection in the midlatitudes.
@@
The scaling CAPE (right hand side of Eq.1) builds a simple linear relation between CAPE and basic environmental parameters of the near-surface and free tropospheric atmospheres, without relying on the profiles of a hypothetically lifted air parcel. The scaling CAPEc, estimated using MSEs and MSE at the time of annual maximum MSEs, effectively captures the spatial distribution of CAPEc (pattern correlation coefficient = 0.98; Fig. 2 and ), with values closely aligned along the one-to-one line (Fig. 2). We assume that MSE remains relatively constant over short time periods, which is consistent with previous studies suggesting that the free troposphere experiences minimal variations during the heatwave evolution (60, 45) or convection build-up periods (55, 56). Consequently, based on Eq.1, CAPE and MSEs are linearly proportional and expected to reach their peak intensities simultaneously. The frequent concurrence of extreme near-surface moist heat and convective instability suggests a shared governing process that drives the accumulation of both phenomena over midlatitude land, thereby the factor determining the maximum MSEs simultaneously sets the maximum CAPE. We proceed to provide a theoretical prediction for both maximum moist heat and maximum potential convection.
@@
The foundation of our framework is grounded in the stored-energy nature of severe convection over midlatitude continents (46, 47, 48, 49, 50, 52, 53, 54), primarily attributed to a preexisting stable lower free troposphere (61, 62, 63, 15, 64, 65, 55, 56, 66). For instance, as shown in a sample sounding from a week prior to the annual maximum surface moist heat in 2004 over the central United States (temperature profiles in Fig. 3), the low-level atmospheric stability typically creates a strong layer of negative buoyancy (i.e., < 0 and CIN > 0) that inhibits convective initiation, even when the free troposphere may already be unstable (i.e., > 0 and CAPE > 0). Therefore, assuming sufficient heat sources, the near-surface atmosphere can continue to warm or moisten until the negative buoyancy layer is eroded, enabling the initiation of deep free convection to release CAPE and terminate the surface heat through precipitation followed by evaporative cooling. In the case study presented, near-surface heat progressively increased over the following days, accompanied by a rise in CAPE. Meanwhile, the free troposphere remained relatively constant until the time of maximum surface moist heat, when convective inhibition was nearly eliminated (temperature profiles in Fig. 3). At the critical point, when both the MSEs and CAPE reach their maximum values, the local atmospheric environment is characterized by CIN 0 and CAPEc 0 (temperature profiles in Fig. 3), consistent with the results presented above (Fig. S2 and 2).
Therefore, the regime of near-surface moist heat accumulation and convection buildup is constrained by CIN 0. Since CIN essentially reflects the presence of negative buoyancy layers within the lower atmosphere, whether and when CIN approaches 0 is governed by the evolution of the most negative buoyancy () throughout the low levels (55, 56). Therefore, CIN 0 is effectively equivalent to 0. The latter is more general and practical, as it also captures cases where CIN is 0 by definition due to the absence of a defined level of free convection (LFC) and equilibrium level (EL), with the entire atmospheric column characterized by negative buoyancy for the near-surface lifted air parcel.
As parcel buoyancy is proportional to the (virtual) temperature difference between the lifted air parcel and its environment, we further utilize the conservation of the parcel’s MSE to derive an alternative formulation of temperature perturbations entirely as a function of environmental static energies (Eq.8, detailed in ). This also provides an MSE perspective to visualize the profiles of temperature perturbations and buoyancy by comparing the parcel’s initial MSE (i.e., ) with the environmental moist static energy profile, which is for or for , where represents the lifted condensation level (Eq.9, detailed in ). As is usually above the (55, 56), based on Eq.9, 0 is equivalent to
@@
The right-hand side of Eqs.3 and 4 defines the maximum potential intensities of near-surface moist heat (MSEs) and convection (CAPE), respectively. These maximum potential intensities are achievable under conditions where surface heat sources are sufficient and external lifting forces are absent, requiring to approach zero to enable free convection, which is commonly observed in midlatitude continents (62, 67, 68, 55, 56). Otherwise, the actual intensities are likely to be lower than their maximum potential values.
@@
We first present the results of the sounding evolutions for the case discussed above, analyzed within the developed MSE framework. As a validation of Eq.9, the MSE profiles (Fig. 3 and ) recover well the positive and negative buoyancy layers and their temporal evolutions. Compared to temperature profiles, the MSE profiles more clearly delineate the most negative buoyancy layer prior to the annual maximum moist heat, identified by a distinct MSE∗ inversion above the LCL (Fig. 3). The preexisting MSE resides at the peak of the MSE∗ inversion near 800 hPa (MSE profiles in Fig. 3), and eventually aligns with the LCL (MSE profiles in Fig. 3). In this case, the final MSE effectively sets the maximum MSEs achieved at the time of the annual maximum moist heat, when (and hence CINc) approaches 0 and CAPEc is maximized. Relative to the significant changes in MSEs during this period, the preexisting MSE intensities only slightly over time, which suggests the possibility of predicting the maximum potential intensity of MSEs (and consequently CAPEc) based on the initial MSE, although adjustments are needed over time as MSE evolves.
@@
To evaluate the generality of the above results, we extend the analysis by presenting composite time series of key proxies derived from a subset of cases over midlatitude land. Specifically, we focus on instances of concurrent annual maxima in MSEs and CAPE for each year during 1980–2022 (44.5% of all cases; Fig. S2). To best visualize the features aligning with our theoretical framework - where MSEs and CAPE are maximized as preexisting CIN is eroded to zero - we further select cases based on the following criteria: CAPEc > 500 J kg-1, CIN 0, and a maximum CIN of at least 50 J kg-1 within seven days preceding the critical time (detailed in ). These selection criteria yield 2502 cases per year (27% of concurrent cases), for which we plot composite 14-day time series centered on the time of maximum MSEs (Fig. 3–). Following the maximum potential for moist heat defined by Eq.3, MSEs increases over time and ultimately converges toward MSE at the peak time (Fig. 3). The height of MSE is in general between 750–800 hPa, and decreases gradually with time to approach (Fig. S3). Notably, MSE fails to predict the maximum MSEs as discussed above, as MSEs exceeds MSE as early as seven days before the peak, with the difference rapidly amplifying over time (Fig. 3). This growing discrepancy between MSEs and MSE reflects the progressive accumulation of CAPE leading up to the peak, as recovered by the scaling CAPE following Eq.1 and constrained by the maximum potential for convection defined in Eq.4 (Fig. 3). Fundamentally, the composite temporal evolutions of MSE and CAPE are linked to the removal of CIN or (Fig. 3), supporting our basic hypothesis formulated in Eq.2. These results are broadly consistent when we perform composite analyses based on all cases of concurrent moist heat and convection over midlatitude land (detailed in ; Fig. S4), though the maximum potential intensities of MSEs and CAPEc are not necessarily achieved (Fig. S4–). This suggests the potential influence of external forcing mechanisms of lift (69, 70, 71), which may facilitate the initiation of convection in cases with non-zero CINc (Fig. S4). Similarly, the composite results reveal relatively steady time series for MSE and MSE (Fig. 3), as well as for the potential CAPE (Fig. 3), highlighting the potential to predict the maximum intensities of MSEs and CAPE based on preexisting free tropospheric properties.
@@
Lastly, to comprehensively validate the maximum potential intensity framework (i.e., Eqs.3 and 4), we compare the observed annual maximum moist heat (MSEs) and the associated critical convection (CAPEc) with their potential maxima derived for all cases over midlatitude land during 1980–2022 from the ERA5 reanalysis data (Fig. 4). Here, MSE and MSE are calculated at the time of annual maximum MSEs. Overall, the maximum potential MSEs (Fig. 4) and the maximum potential CAPEc (Fig. 4) capture the spatial patterns of the observed maximum MSEs (Fig. 1) and CAPEc (Fig. 2) over midlatitude land, with observed maxima slightly smaller than their potential counterparts. The pattern correlation coefficient is 0.94 between observed and potential MSEs maxima, and 0.82 between observed and potential CAPEc. The joint histograms of all cases further confirm that the maximum potential intensities serve as a tight upper limit for the observed maximum MSEs and CAPEc (Fig. 4 and ). For the maximum MSEs, the potential MSEs is achievable for many cases, with others closely aligned with the one-to-one line but slightly below (Fig. 4). Translating (potential) MSEs into (potential) WBTs (as detailed in Materials and Methods) yields consistent results (Fig. 4), highlighting the applicability of our theory for directly predicting maximum WBTs. Similarly, the potential CAPEc serves as a strong constraint on the observed CAPEc (Fig. 4), although the actual CAPEc exhibits greater variability for a given potential CAPEc, partly due to the variations in MSEs.

### v4

@@ Abstract
Extreme near-surface moist heat and severe convective storms are among the leading causes of weather-related damages worldwide. Here, we show that episodes of extreme moist heat and severe convection frequently co-occur across midlatitude land regions, and develop a theoretical framework that links their maximum potential intensities to preexisting low-level energy inversions. By accounting for the stored-energy nature of midlatitude severe convection, where moist heat and atmospheric instability accumulate before convection initiates, our work advances the understanding of convective constraints on extreme heat events. The theory identifies low-level inversions as a critical factor shaping compound extreme heat and convective weather risks, and offers a pathway for improving the modeling and future projection of these events.
keywords:
atmospheric stability moist heat severe convection convective inhibition low-level energy inversionThis manuscript was compiled on August 6, 2025
@@
Moist heatwaves, which combine the effects of high temperature and humidity, pose significant risks to public health and societal outcomes (1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15). While previous research has primarily focused on the dynamics and changes of moist heat in tropical and subtropical regions (16, 17, 18, 19, 20, 21, 22), the drivers and characteristics of extreme moist heat over midlatitude continents remain poorly understood. As climate change is driving severe moist heat risks poleward into the midlatitudes (23, 24, 25, 26, 27, 28, 29, 30), advancing our theoretical understanding of the limits of maximum moist heat in these regions has become increasingly urgent. Our goal in this study is to determine what sets the maximum moist heat over midlatitude land.
@@
Near-surface wet bulb temperature (WBTs) is a key measure of moist heat stress, as it combines the influence of temperature and humidity (31, 4, 32, 6, 26, 19, 29, 33). By definition, WBT represents the temperature that an air parcel would attain if cooled adiabatically to saturation at constant pressure by evaporating water into it (31, 34, 35). During this adiabatic process, the air parcel conserves its moist static energy (MSE) or the equivalent potential temperature, and thus WBTs can be directly derived assuming the conservation of near-surface MSE (MSEs) (36, 19, 18, 33, 37). Unlike the non-linear WBT formula, MSE is a linear function of temperature, specific humidity, and geopotential height, making it simple to use (18, 38). Accordingly, this study employs MSEs, defined as the sum of sensible heat, latent heat, and geopotential energy at 2 meters above the ground, to measure near-surface moist heat. Overall, MSEs reasonably represents moist heat and, similar to the annual maximum WBTs (Fig. 1), the annual maximum MSEs recovers well the spatial distribution of extreme moist heat over midlatitude land (Fig. 1) and aligns closely with WBTs (Fig. S1; detailed in ). In this work, we use the 3-hourly ERA5 reanalysis data (39, 40) during 1980–2022 for land grids between 35∘N and 75∘N with elevation lower than 1000 m (detailed in ). Our analysis focuses on the historical annual maximum MSEs to explore the theoretical constraint on the maximum moist heat over midlatitude land.
@@
Previous studies have commonly assumed a strong coupling between the near-surface atmosphere and the free troposphere during extreme heat through deep moist convection, such that MSEs is constrained by free troposheric saturated moist static energy (MSE∗) given the state of convective quasi-equilibrium (41, 42, 43, 4, 44, 45, 46, 47, 19, 17, 6). Under the assumption of a moist-neutral atmospheric column, this framework has been widely used to study extreme heat in the tropics (17, 19, 21) or the global mean state (6). With the neutrality assumption, a recent study (48) has applied this framework to investigate extreme dry heat over midlatitude continents, suggesting that MSEs is limited by 500-hPa MSE∗ (i.e., MSEMSE), thereby setting an upper bound for maximum surface air temperature. We first test the validity of this theory by comparing MSEs with MSE during the annual maximum moist heat over midlatitude land. Our analysis reveals that while the theory (and the neutrality assumption) might be more correct in the dry limit (48), it does not hold for extreme moist heat, as MSEs is consistently and significantly greater than MSE in nearly all cases (more than 96% of the midlatitude land grid points; Fig. 1–). Furthermore, the distribution of MSE is largely zonally uniform (Fig. 1), which does not capture the strong zonal variation observed in the MSEs pattern (Fig. 1).
@@
These results indicate a violation of the moist neutrality assumption because, unlike quasi-equilibrium convection in the tropics, severe convection over midlatitude continents often involves a strong accumulation of convective energy (49, 50, 51, 52, 53, 54, 55, 56, 57, 58), which may interact with surface heating in distinct ways. Although this stored-energy nature of midlatitude severe convection has been well recognized within the severe weather community (59, 60), it has not been formulated to constrain either the maximum potential intensity of convection (54, 55, 57, 58) or the maximum surface heat available at the onset of convection (48). Here, building on the stored-energy energy of severe convection and the common assumption that convective precipitation limits surface heat, we propose a new theoretical framework that accounts for the onset of convection, enables surface heat and convective instability to evolve beyond the neutral condition, and ultimately provides a tight constraint on both near-surface moist heat and potential convection over midlatitude land.
In the main text, we present and validate our theoretical framework within current climate using ERA5 reanalysis data. We conclude by discussing the important implications of our results for predicting and understanding the maximum intensities of severe convection and near-surface moist heat in future climates.
@@
We first demonstrate that deep convection (Fig. 2) occurs concurrently with the annual maximum moist heat (Fig. 1), supporting the hypothesis that convective instability acts to terminate heatwaves and constrain maximum near-surface heat (6, 48, 61, 62, 20). The potential intensity of convection is measured by convective available potential energy (CAPE; detailed in ), which is defined as the vertical integral of positive parcel buoyancy through the deep troposphere, with values above 1000 J kg-1 typically deemed sufficient to support strong to severe convection (63). We define CAPE at the time of the annual maximum MSEs as critical CAPE (denoted CAPEc), which measures the potential convection that terminates the annual maximum moist heat. The CAPEc over most midlatitude land regions exceeds 1000 J kg-1, with values ranging from 3000 to 4000 J kg-1 across western Europe extending to Northeast China, and reaching 6000 J kg-1 or higher over eastern China and central North America (Fig. 2). The high values of CAPEc represent substantial convective instability in the atmosphere during extreme moist heat, indicating a significant departure from the neutrality assumption. Meanwhile, critical convective inhibition (denoted CINc; detailed in ), a measure of the vertically integrated energy barrier in the lower free troposphere that suppresses the release of CAPE, is sufficiently low (generally less than 25 J kg-1) to permit the initiation of free convection (Fig. S2).
These high-CAPE, low-CIN environments largely ensure the occurrence of deep convection that terminates extreme near-surface heat, as further evidenced by substantial precipitation following the hottest MSEs day across most midlatitudes regions (Fig. S2). In fact, the deep convection that terminates the maximum moist heat is likely among the most intense convection of the year across many midlatitude land areas, as CAPEc (Fig. 2) aligns closely with the annual maximum CAPE (Fig. S2) and the precipitation following peak MSEs generally ranks in the high percentiles (top 10–20%) of daily precipitation during the year (Fig. S2). This alignment is particularly evident in Europe and central North America, where the difference between CAPEc and the annual maximum CAPE is less than 250 J kg-1 (Fig. S2). In these regions, over 50% of the years during 1980–2022 feature simultaneous annual maxima of CAPE and MSEs (Fig. S2), emphasizing the frequent co-occurrence of extreme moist heat and intense convection in the midlatitudes.
Note that several regions including the land east of the Caspian Sea, south of the Mediterranean Sea, and along the west coast of the United States, exhibit high MSEs maxima (Fig. 1–) but only modest CAPEc (Fig. 2), however, the relatively high CINc (up to 100 J kg-1 or greater) found in these regions (Fig. S2) likely inhibits convective initiation and strongly limits precipitation there (Fig. S2). These patterns are consistent with the weak overlap between the maxima of CAPE and MSEs (Fig. S2), as well as the low annual percentiles of precipitation following peak MSEs (Fig. S2), which indicates a less frequent co-occurrence of extreme moist heat and intense convection, and thus processes other than convective precipitation may be responsible for terminating extreme heat in these regions.
@@
The scaling CAPE (right hand side of Eq.1) builds a simple linear relation between CAPE and basic environmental parameters of the near-surface and free tropospheric atmospheres, without relying on the profiles of a hypothetically lifted air parcel. The scaling CAPEc, estimated using MSEs and MSE at the time of annual maximum MSEs, effectively captures the spatial distribution of CAPEc (pattern correlation coefficient = 0.98; Fig. 2 and ), with values closely aligned along the one-to-one line (Fig. 2). Note that the virtual temperature correction is considered in our calculation of standard CAPE but ignored in our derivation of scaling CAPE for simplicity, but the strong agreement between the two suggests that virtual effects are minor. Since CAPEc is generally high, this is consistent with previous findings that virtual temperature correction is negligible for large CAPE values (64).
If MSE does not change much during the heatwave evolution and convection build-up periods (consistent with results found by previous studies (65, 48, 57, 58) and supported by Fig. 3 discussed later), then from Eq.1, CAPE and MSEs are expected to reach their peak intensities roughly at the same time. The frequent concurrence of extreme near-surface moist heat and convective instability suggests a shared governing process that drives the accumulation of both phenomena over midlatitude land, thereby the factor determining the maximum MSEs simultaneously sets the maximum CAPE. We proceed to provide a theoretical prediction for both maximum moist heat and maximum potential convection.
@@
The foundation of our framework is grounded in the stored-energy nature of severe convection over midlatitude continents (49, 59, 50, 51, 52, 54, 55, 56), primarily attributed to a preexisting stable lower free troposphere (66, 67, 68, 18, 69, 70, 57, 58, 71). For instance, as shown in a sample sounding from 5 days prior to the annual maximum surface moist heat in 2004 over the central United States (temperature profiles in Fig. 3), the low-level atmospheric stability typically creates a strong layer of negative buoyancy (i.e., < 0 and CIN > 0; initial CIN=180 J kg-1 for the case) that inhibits convective initiation, even when the free troposphere may already be unstable (i.e., > 0 and CAPE > 0; initial CAPE=672 J kg-1 for the case). Therefore, assuming sufficient heating sources, the near-surface atmosphere can continue to warm or moisten until the negative buoyancy layer is eroded, enabling the initiation of deep free convection to release CAPE and terminate the surface heat through precipitation followed by evaporative cooling. In the case study presented, near-surface heat progressively increased over the following days, accompanied by a rise in CAPE. Meanwhile, the free troposphere remained relatively constant until the time of maximum surface moist heat, when convective inhibition was nearly eliminated (temperature profiles in Fig. 3). At the critical point, when both the MSEs and CAPE reach their maximum values, the local atmospheric environment is characterized by CIN 0 and CAPEc 0 (temperature profiles in Fig. 3; final CIN=9 J kg-1 and CAPE=6753 J kg-1 for the case), consistent with the results presented above (Fig. S2 and 2).
Therefore, the regime of near-surface moist heat accumulation and convection buildup is constrained by CIN 0. Since CIN essentially reflects the presence of negative buoyancy layers within the lower atmosphere, whether and when CIN approaches 0 is governed by the evolution of the most negative buoyancy () throughout the low levels (57, 58). Therefore, CIN 0 is effectively equivalent to 0. The latter is more general and practical, as it also captures cases where CIN is 0 by definition due to the absence of a defined level of free convection (LFC) and equilibrium level (EL), with the entire atmospheric column characterized by negative buoyancy for the near-surface lifted air parcel.
As parcel buoyancy is proportional to the (virtual) temperature difference between the lifted air parcel and its environment, we further utilize the conservation of the parcel’s MSE to derive an alternative formulation of temperature perturbations entirely as a function of environmental static energies (Eq.8, detailed in ). This also provides an MSE perspective to visualize the profiles of temperature perturbations and buoyancy by comparing the parcel’s initial MSE (i.e., ) with the environmental moist static energy profile, which is for or for , where represents the lifted condensation level (Eq.9, detailed in ). As is usually above the (57, 58), based on Eq.9, 0 is equivalent to
@@
The right-hand side of Eqs.3 and 4 defines the maximum potential intensities of near-surface moist heat (MSEs) and convection (CAPE), respectively. These maximum potential intensities are achievable under conditions where surface heat sources are sufficient and external lifting forces are absent, requiring to approach zero to enable free convection, which is commonly observed in midlatitude continents (67, 72, 73, 57, 58). Otherwise, the actual intensities are likely to be lower than their maximum potential values.
@@
We first present the results of the sounding evolutions for the case discussed above, analyzed within the developed MSE framework. Those profiles, taken from the ERA5 reanalysis data, closely align with radiosonde observations (Fig. S3). As a validation of Eq.9, the MSE profiles (Fig. 3 and ) recover well the positive and negative buoyancy layers and their temporal evolutions. Compared to temperature profiles, the MSE profiles more clearly delineate the most negative buoyancy layer prior to the annual maximum moist heat, identified by a distinct MSE∗ inversion above the LCL (Fig. 3). The preexisting MSE resides at the peak of the MSE∗ inversion near 800 hPa (MSE profiles in Fig. 3), and eventually aligns with the LCL (MSE profiles in Fig. 3). In this case, the final MSE effectively sets a constraint that is only slightly greater than the maximum MSEs achieved at the time of the annual maximum moist heat, when (and hence CINc) approaches nearly 0 and CAPEc is maximized (CINc=9 J kg-1 and CAPEc=6753 J kg-1). The small departure of peak MSEs form the final MSE is consistent with the small residual of CINc. Relative to the significant changes in MSEs during this period, the preexisting MSE intensities weakly over time, which suggests the possibility of predicting the maximum potential intensity of MSEs (and consequently CAPEc) based on the initial MSE, although adjustments are needed over time as MSE evolves.
@@
To evaluate the generality of the above results, we extend the analysis by presenting composite time series of key proxies derived from a subset of cases over midlatitude land. Specifically, we focus on instances of concurrent annual maxima in MSEs and CAPE for each year during 1980–2022 (44.5% of all cases; Fig. S2). To best visualize the features aligning with our theoretical framework - where MSEs and CAPE are maximized as preexisting CIN is eroded to zero - we further select cases based on the following criteria: CAPEc > 500 J kg-1, CIN 0, and a maximum CIN of at least 50 J kg-1 within seven days preceding the critical time (detailed in ). These selection criteria yield 2502 cases per year (27% of concurrent cases), for which we plot composite 14-day time series centered on the time of maximum MSEs (Fig. 3–). Following the maximum potential for moist heat defined by Eq.3, MSEs increases over time and ultimately converges toward MSE at the peak time (Fig. 3). The height of MSE is in general between 750–800 hPa, and decreases gradually with time to approach (Fig. S4). Notably, MSE fails to predict the maximum MSEs as discussed above, as MSEs exceeds MSE as early as seven days before the peak, with the difference rapidly amplifying over time (Fig. 3). This growing discrepancy between MSEs and MSE reflects the progressive accumulation of CAPE leading up to the peak, as recovered by the scaling CAPE following Eq.1 and constrained by the maximum potential convection defined in Eq.4 (Fig. 3). The peak CAPE on the hottest MSEs day, followed by its decrease, indicates the onset of convection that consumes convective instability and terminates extreme heat, which is further reflected by daily precipitation that also peaks on the hottest day and remains relatively high over the following days (Fig. S4). Fundamentally, the composite temporal evolutions of MSE and CAPE are linked to the removal of CIN or (Fig. 3), supporting our basic hypothesis formulated in Eq.2. These results are broadly consistent when we perform composite analyses based on all cases of concurrent moist heat and convection over midlatitude land (detailed in ; Fig. S5), though the maximum potential intensities of MSEs and CAPEc are not necessarily achieved (Fig. S5–). This suggests the potential influence of external forcing mechanisms of lift (74, 75, 76), which may facilitate the initiation of convection in cases with non-zero CINc (Fig. S5). Similarly, the composite results reveal relatively steady time series for MSE and MSE (Fig. 3), as well as for the potential CAPE (Fig. 3), highlighting the potential to predict the maximum intensities of MSEs and CAPE based on preexisting free tropospheric properties.
@@
Lastly, to comprehensively validate the maximum potential intensity framework (i.e., Eqs.3 and 4), we compare the observed annual maximum moist heat (MSEs) and the associated critical convection (CAPEc) with their potential maxima derived for all cases over midlatitude land during 1980–2022 from the ERA5 reanalysis data (Fig. 4). Here, MSE and MSE are calculated at the time of annual maximum MSEs. Overall, the maximum potential MSEs (Fig. 4) and the maximum potential CAPEc (Fig. 4) capture the spatial patterns of the observed maximum MSEs (Fig. 1) and CAPEc (Fig. 2) over midlatitude land, with observed maxima slightly smaller than their potential counterparts. The pattern correlation coefficient is 0.94 between observed and potential MSEs maxima, and 0.82 between observed and potential CAPEc. The joint histograms of all cases further confirm that the maximum potential intensities serve as a tight upper limit for the observed maximum MSEs and CAPEc (Fig. 4 and ). For the maximum MSEs, the potential MSEs is achievable for many cases, with others closely aligned with the one-to-one line but slightly below (Fig. 4). Translating (potential) MSEs into (potential) WBTs (as detailed in Materials and Methods) yields consistent results (Fig. 4), highlighting the applicability of our theory for directly predicting maximum WBTs. Similarly, the potential CAPEc serves as a strong constraint on the observed CAPEc (Fig. 4), although the actual CAPEc exhibits greater variability for a given potential CAPEc, partly due to the variations in MSEs. It is worth noting the relatively large discrepancy over the land east of the Caspian Sea, south of the Mediterranean Sea, and along the west coast of the United States (Fig. 4– versus Fig. 1 and 2), where the residual energy barrier (i.e., the difference between MSE and MSEs at the peak time) is high, consistent with the high CINc discussed above. This again suggests that processes other than convection are responsible for terminating extreme heat events before surface heating and convection can approach their potential intensities in regions where convective inhibition persists.